what volume of 5.00X10^-3 M HNO3 is needed to titrate 80.00 mL of 5.00X 10^-3 M Ca(OH)2 to the equivalence point
Answer #1

The reaction is the following:

2HNO3 + Ca(OH)2 -----------> Ca(NO3)2 + 2H2O

With this reaction, we see that it takes 2 moles of HNO3 to react with 1 mole of Ca(OH)2 so:

mol A = mol B and in this case:

mole HNO3 = 2 moles

mole Ca(OH)2 = 1 mole

mole HNO3 / mole Ca(OH)2 = 2/1

mole HNO3 = 2 mole Ca(OH)2

If mole = M * V then:

MaVa = 2MbVb

0.005 * Va = 2 * 0.005 * 80

Va = 160 mL